### v4

@@ 1.1 Methods
One such instance is in our improved bound for order-finding and exponentiation in a semigroup (Lem. 3.1). The previous proof [BKLM01] (still state of the art 23 years later) used a then-novel and clever “double-barrelled” recursive approach to compute these in $\mathsf{FOLL}$. In contrast, our proof uses standard repeated doubling, noting that it can be done in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$, recovering their result with standard tools and reducing the depth; in fact, from our proof we get $\mathsf{quasiAC^{0}}$ circuits of depth 2, which is clearly optimal from the perspective of depth. We use this improved bound on order-finding to improve the complexity of isomorphism testing of Abelian groups (Thm. 4.1), simple groups (Cor. 3.3), and groups of almost all orders (Thm. 5.1).
@@
The first key is Thm. D, putting Membership for quasigroups into $\mathsf{NTIME}(\mathrm{polylog}(n))$ (more precisely, into $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$). To do this, we replace the use of a cube generating sequences from [CTW13] with something that is nearly as good for the purposes of MGS: we extend the Babai–Szemerédi Reachability Lemma [BS84, Thm. 3.1] from groups (its original setting) to quasigroups in order to obtain what we call cube-like generating sequences, which also give us short straight-line programs. Division in quasigroups is somewhat nuanced, e.g., despite the fact that for any $a,b$, there exists a unique $x$ such that $ax=b$, this does not necessarily mean that there is an element “$a^{-1}$” such that $x=a^{-1}b$, because of the lack of associativity. Our proof is thus a careful adaptation of the technique of Babai & Szemerédi, with a few quasigroup twists that result in a slightly worse, but still sufficient, bound.
@@ 1.2 Prior work
The best known runtime bound for GpI is $n^{(1/4)\log_{p}(n)+O(1)}$-time due to Rosembaum [Ros13] (see [LGR16, Sec. 2.2]), though this tells us little about parallel complexity. In addition to Tarjan’s result mentioned above [Mil78], Lipton, Snyder, & Zalcstein [LSZ77] independently observed that if a group is $d$-generated, then we can decide isomorphism by considering all possible $d$-element subsets. This is the generator enumeration procedure. Using the fact that every group admits a generating set of size $\leq\log_{p}(n)$ (where $p$ is the smallest prime dividing $n$), Tarjan obtained a bound of $n^{\log_{p}(n)+O(1)}$-time for Group Isomorphism, while Lipton, Snyder, & Zalcstein [LSZ77] obtained a stronger bound of $\textsf{DSPACE}(\log^{2}n)$. Miller [Mil78] extended Tarjan’s observation to the setting of quasigroups. There has been subsequent work on improving the parallel complexity of generator enumeration for quasigroups, resulting in bounds of $\exists^{\log^{2}n}\mathsf{AC}^{1}$ ($\mathsf{AC}^{1}$ circuits that additionally receive $O(\log^{2}n)$ non-deterministic bits, denoted by other authors as $\beta_{2}\mathsf{AC}^{1}$) due to Wolf [Wol94], [^12] See footnote 6. $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ due to Wagner [Wag10], and $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ due to Chattopadhyay, Torán, & Wagner [CTW13]. In the special case of groups, generator enumeration is also known to belong to $\exists^{\log^{2}n}\textsf{SC}^{2}$ [Tan13]. There has been considerable work on polynomial-time, isomorphism tests for several families of groups, as well as more recent work on NC isomorphism tests—we refer to recent works [GQ17, DW22, GL23] for a survey. We are not aware of work on isomorphism testing for specific families of quasigroups that are not groups.
@@
As every (quasi)group has a generating set of size $\leq\lceil\log n\rceil$, MGS admits an $n^{\log(n)+O(1)}$-time solution for (quasi)groups. In the case of groups, Arvind & Torán [AT06] improved the complexity to $\textsf{DSPACE}(\log^{2}n)$. They also gave a polynomial-time algorithm in the special case of nilpotent groups. Tang further improved the general bound for MGS for groups to $\exists^{\log^{2}n}\textsf{SC}^{2}$ [Tan13]. We observe that Wolf’s technique for placing Quasigroup Isomorphism into $\mathsf{DSPACE}(\log^{2}n)$ also suffices to get MGS for Quasigroups into the same class. Very recently, Das & Thakkar [DT23] improved the algorithmic upper bound for MGS in the setting of groups to $n^{(1/4)\log(n)+O(1)}$. A month later, Lucchini & Thakkar [LT24] placed MGS for groups into P. Prior to [LT24], MGS for Groups was considered comparable to Group Isomorphism in terms of difficulty. Our $\mathsf{AC}^{1}(\mathsf{L})$ bound (Thm. C) further closes the gap between Membership Testing in groups and MGS for Groups, and in particular suggests that MGS is of comparable difficulty to Membership for groups rather than GpI. Note that Membership for groups has long been known to belong to $\mathsf{L}$ [BM91, Rei08].
@@ 2.1 Algebra and Combinatorics
Graph Theory. A strongly regular graph with parameters $(n,k,\lambda,\mu)$ is a simple, undirected $k$-regular, $n$-vertex graph $G(V,E)$ where any two adjacent vertices share $\lambda$ neighbors, and any two non-adjacent vertices share $\mu$ neighbors. The complement of a strongly regular graph is also strongly regular, with parameters $(n,n-k-1,n-2-2k+\mu,n-2k+\lambda)$.
@@
Quasigroups and Latin squares. A quasigroup consists of a set $G$ and a binary operation $\star:G\times G\to G$ satisfying the following. For every $a,b\in G$, there exist unique $x,y$ such that $a\star x=b$ and $y\star a=b$. We write $x=a\backslash b$ and $y=b/a$, i. e., $a\star(a\backslash b)=b$ and $(b/a)\star a=b$. When the multiplication operation is understood, we simply write $ax$ for $a\star x$. A sub-quasigroup of a quasigroup is a subset that itself is a quasigroup. This means it is closed under the multiplication as well as under left and right quotients. Given $X\subseteq G$, the sub-quasigroup generated by $X$ is denoted as $\langle X\rangle$. It is the smallest sub-quasigroup containing $X$.
@@
We say that $S$ is a cube generating sequence if each element $g$ in the quasigroup can be written as $g=P(s_{0}s_{1}^{e_{1}}\cdots s_{k}^{e_{k}})$, for $e_{1},\ldots,e_{k}\in\{0,1\}$. Here, $s_{i}^{0}$ indicates that $s_{i}$ is not being considered in the product. For every parenthesization, every quasigroup is known to admit a cube generating sequence of size $O(\log n)$ [CTW13].
@@
Albert showed that a quasigroup $Q$ is isotopic to a group $G$ if and only if $Q$ is isomorphic to $G$. In general, isotopic quasigroups need not be isomorphic [Alb43].
@@
A Latin square $L$ can equivalently be viewed as a set of triples $\{(i,j,L_{ij}):i,j\in[n]\}\subseteq[n]\times[n]\times[n]$. Given a set of triples $S\subseteq[n]\times[n]\times[n]$, the Latin square property can equivalently be rephrased as: every $i\in[n]$ appears as the first—resp. second, resp. third—coordinate of some triple in $S$, and no two triples in $S$ agree in more than one coordinate. From this perspective, an additional potential symmetry of Latin squares emerges: two Latin squares $L_{1},L_{2}$ are parastrophic [^13] This terminology is borrowed from the quasigroup literature. In the Latin square literature this is sometimes referred to as “conjugate”, but we find the argument in Keedwell and Denes [KD15, pp. 15–16] to use the quasigroup nomenclature even in the setting of Latin squares compelling. if there is a permutation $\pi\in S_{3}$ such that, when viewed as sets of triples, we have $L_{2}=\{(x_{1^{\pi}},x_{2^{\pi}},x_{3^{\pi}}):(x_{1},x_{2},x_{3})\in L_{1}\}$; the induced map $L_{1}\to L_{2}$ is called a parastrophy. A main class isomorphism of Latin squares is the composition of a parastrophy and an isotopy; if there exists a main class isomorphism $L_{1}\to L_{2}$, we say they are main class isomorphic. [^14] Miller [Mil78], and then Levet [Lev23] following Miller, referred to this as “main class isotopy”; we have since found several modern textbooks on quasigroups and Latin squares that refer to this notion as one of “main class isomorphism”, “paratopy”, or “isostrophy” (sic). Keedwell and Denes [KD15, pp. 15–16] have a nice discussion of the terminology, as well as the history of its usage.
@@
For a given Latin square $L$ of order $n$, we associate a Latin square graph $G(L)$ that has $n^{2}$ vertices; one for each triple $(a,b,c)$ that satisfies $ab=c$. Two vertices $(a,b,c)$ and $(x,y,z)$ are adjacent in $G(L)$ precisely if $a=x$ or $b=y$ or $c=z$. Miller showed that two Latin square graphs $G_{1}$ and $G_{2}$ are isomorphic if and only if the corresponding Latin squares, $L_{1}$ and $L_{2}$, are main class isomorphic [Mil78].
@@
A Latin square graph on $n^{2}$ vertices is a strongly regular graph with parameters $(n^{2},3(n-1),n,6)$. Conversely, a strongly regular graph with these same parameters $(n^{2},3(n-1),n,6)$ is called a pseudo-Latin square graph. Bruck showed that for $n>23$, a pseudo-Latin square graph is a Latin square graph [Bru63].
@@
Group Theory. For a standard reference, see [Rob82]. All groups are assumed to be finite. For a group $G$, $d(G)$ denotes the minimum size of a generating set for $G$. The Frattini subgroup $\Phi(G)$ is the set of non-generators of $G$. If $G$ is a $p$-group, the Burnside Basis Theorem (see [Rob82, Theorem 5.3.2]) provides that (i) $G=G^{p}[G,G]$, (ii) $G/\Phi(G)\cong(\mathbb{Z}/p\mathbb{Z})^{d(G)}$, and (iii) if $S$ generates $(\mathbb{Z}/p\mathbb{Z})^{d(G)}$, then any lift of $S$ generates $G$. A chief series of $G$ is an ascending chain $(N_{i})_{i=0}^{k}$ of normal subgroups of $G$, where $N_{0}=1$, $N_{k}=G$, and each $N_{i+1}/N_{i}$ ($i=0,\ldots,k-1$) is minimal normal in $G/N_{i}$. For $g,h\in G$, the commutator $[g,h]:=ghg^{-1}h^{-1}$. The commutator subgroup $[G,G]=\langle\{[g,h]:g,h\in G\}\rangle$.
@@
Designs. Let $t\leq k\leq v$ and $\lambda$ be positive integers. A $(t,k,\lambda,v)$ design is an incidence structure $\mathcal{D}=(X,\mathcal{B},I)$, where $X$ is a set of $v$ points, $\mathcal{B}$ is a subset of $\binom{X}{k}$—whose elements are referred to as blocks—and such that each $t$-element subset of $X$ belongs to exactly $\lambda$ blocks. Now $I$ is the point-block incidence matrix, where $I_{x,B}=1$ precisely if the point $x$ belongs to the block $B$.
@@
For a design $\mathcal{D}=(X,\mathcal{B},I)$, we may define a block intersection graph (also known as a line graph) $G(V,E)$, where $V(G)=\mathcal{B}$ and two blocks $B_{1},B_{2}$ are adjacent in $G$ if and only if $B_{1}\cap B_{2}\neq\emptyset$. In the case of a Steiner $2$-design, the block-intersection graph is strongly regular. For Steiner triple systems, the block-intersection graphs are strongly regular with parameters $(n(n-1)/6,3(n-3)/2,(n+3)/2,9)$. Conversely, strongly regular graphs with the parameters $(n(n-1)/6,3(n-3)/2,(n+3)/2,9)$ are referred to as pseudo-STS graphs. Bose showed that pseudo-STS graphs graphs with strictly more than $67$ vertices are in fact STS graphs [Bos63].
Algorithmic Problems. A multiplication table of a magma $G=\{g_{1},\dots,g_{n}\}$ of order $n$ is an array $M$ of length $n^{2}$ where each entry $M[i+(j-1)n]$ for $i,j\in\{1,\dots,n\}$ contains the binary representation of $k$ such that $g_{i}g_{j}=g_{k}$. In the following all magmas are given as their multiplication tables.
@@
We will be primarily interested in Minimum Generating Set and Membership in the setting of (quasi)groups, with a few excursions to semigroups and magmas. Note that Membership for groups is known to be in $\mathsf{L}\cap\mathsf{quasiAC^{0}}$ [BM91, Rei08, Fle22]. Here the containment in $\mathsf{L}$ follows from the deep result by Reingold [Rei08] that symmetric logspace (nondeterministic logspace where each transition is also allowed to be applied backward) coincides with $\mathsf{L}$.

### v5

@@ 1.1 Methods
One such instance is in our improved bound for order-finding and exponentiation in a semigroup (Lem. 3.1). The previous proof [10] (still state of the art 23 years later) used a then-novel and clever “double-barrelled” recursive approach to compute these in $\mathsf{FOLL}$. In contrast, our proof uses standard repeated doubling, noting that it can be done in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{FOLL}\cap\mathsf{quasiAC^{0}}$, recovering their result with standard tools and reducing the depth; in fact, from our proof we get $\mathsf{quasiAC^{0}}$ circuits of depth 2, which is clearly optimal from the perspective of depth. We use this improved bound on order-finding to improve the complexity of isomorphism testing of Abelian groups (Thm. 4.1), simple groups (Cor. 3.5), and groups of almost all orders (Thm. 5.1).
@@
The first key is Thm. 1.7, putting Membership for quasigroups into $\mathsf{NTIME}(\mathrm{polylog}(n))$ (more precisely, into $\exists^{\log^{2}n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$). To do this, we replace the use of a cube generating sequences from [17] with something that is nearly as good for the purposes of MGS: we extend the Babai–Szemerédi Reachability Lemma [7, Thm. 3.1] from groups (its original setting) to quasigroups in order to obtain what we call cube-like generating sequences, which also give us short straight-line programs. Division in quasigroups is somewhat nuanced, e.g., despite the fact that for any $a,b$, there exists a unique $x$ such that $ax=b$, this does not necessarily mean that there is an element $a^{-1}$” such that $x=a^{-1}b$, because of the lack of associativity. Our proof is thus a careful adaptation of the technique of Babai & Szemerédi, with a few quasigroup twists that result in a slightly worse, but still sufficient, bound.
@@ 1.2 Prior work
The best known runtime bound for GpI is $n^{(1/4)\log_{p}(n)+O(1)}$-time (where $p$ is the smallest prime dividing $n$) due to Rosenbaum [49] and Luks [41] (see [36, Sec. 2.2]), though this tells us little about parallel complexity. In addition to Tarjan’s result mentioned above [43], Lipton, Snyder, & Zalcstein [38] independently observed that if a group is $d$-generated, then we can decide isomorphism by considering all possible $d$-element subsets. This is the generator enumeration procedure. Using the fact that every group admits a generating set of size $\leq\log_{p}(n)$ (where $p$ is the smallest prime dividing $n$), Tarjan obtained a bound of $n^{\log_{p}(n)+O(1)}$-time for Group Isomorphism, while Lipton, Snyder, & Zalcstein [38] obtained a stronger bound of $\mathsf{DSPACE}(\log^{2}n)$. Miller [43] extended Tarjan’s observation to the setting of quasigroups. There has been subsequent work on improving the parallel complexity of generator enumeration for quasigroups, resulting in bounds of $\exists^{\log^{2}n}\mathsf{AC}^{1}$ ($\mathsf{AC}^{1}$ circuits that additionally receive $O(\log^{2}n)$ non-deterministic bits, denoted by other authors as $\beta_{2}\mathsf{AC}^{1}$) due to Wolf [59], [^12] See footnote 6. $\exists^{\log^{2}n}\mathsf{SAC}^{1}$ due to Wagner [58], and $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ due to Chattopadhyay, Torán, & Wagner [17]. In the special case of groups, generator enumeration is also known to belong to $\exists^{\log^{2}n}\mathsf{SC}^{2}$ [53]. There has been considerable work on polynomial-time, isomorphism tests for several families of groups, as well as more recent work on $\mathsf{NC}$ isomorphism tests—we refer to recent works [27, 22, 26] for a survey. We are not aware of work on isomorphism testing for specific families of quasigroups that are not groups.
@@
As every (quasi)group has a generating set of size $\leq\lceil\log n\rceil$, MGS admits an $n^{\log(n)+O(1)}$-time solution for (quasi)groups. In the case of groups, Arvind & Torán [3] improved the complexity to $\mathsf{DSPACE}(\log^{2}n)$. They also gave a polynomial-time algorithm in the special case of nilpotent groups. Tang further improved the general bound for MGS for groups to $\exists^{\log^{2}n}\mathsf{SC}^{2}$ [53]. We observe that Wolf’s technique for placing Quasigroup Isomorphism into $\mathsf{DSPACE}(\log^{2}n)$ also suffices to get MGS for Quasigroups into the same class. Very recently, Das & Thakkar [20] improved the algorithmic upper bound for MGS in the setting of groups to $n^{(1/4)\log(n)+O(1)}$. A month later, Lucchini & Thakkar [40] placed MGS for groups into $\mathsf{P}$. Prior to [40], MGS for Groups was considered comparable to Group Isomorphism in terms of difficulty. Our $\mathsf{AC}^{1}(\mathsf{L})$ bound (Thm. 1.5) further closes the gap between Membership Testing in groups and MGS for Groups, and in particular suggests that MGS is of comparable difficulty to Membership for groups rather than GpI. Note that Membership for groups has long been known to belong to $\mathsf{L}$ [11, 47].
@@ 2.1 Algebra and Combinatorics
Graph Theory.
A strongly regular graph with parameters $(n,k,\lambda,\mu)$ is a simple, undirected $k$-regular, $n$-vertex graph $G(V,E)$ where any two adjacent vertices share $\lambda$ neighbors, and any two non-adjacent vertices share $\mu$ neighbors. The complement of a strongly regular graph is also strongly regular, with parameters $(n,n-k-1,n-2-2k+\mu,n-2k+\lambda)$.
@@
Quasigroups and Latin squares.
[…]
@@
We say that $S$ is a cube generating sequence if each element $g$ in the quasigroup can be written as $g=P(s_{0}s_{1}^{e_{1}}\cdots s_{k}^{e_{k}})$, for $e_{1},\ldots,e_{k}\in\{0,1\}$. Here, $s_{i}^{0}$ indicates that $s_{i}$ is not being considered in the product. For every parenthesization, every quasigroup is known to admit a cube generating sequence of size $O(\log n)$ [17].
@@
Albert showed that a quasigroup $Q$ is isotopic to a group $G$ if and only if $Q$ is isomorphic to $G$. In general, isotopic quasigroups need not be isomorphic [1].
@@
A Latin square $L$ can equivalently be viewed as a set of triples $\{(i,j,L_{ij}):i,j\in[n]\}\subseteq[n]\times[n]\times[n]$. Given a set of triples $S\subseteq[n]\times[n]\times[n]$, the Latin square property can equivalently be rephrased as: every $i\in[n]$ appears as the first—resp. second, resp. third—coordinate of some triple in $S$, and no two triples in $S$ agree in more than one coordinate. From this perspective, an additional potential symmetry of Latin squares emerges: two Latin squares $L_{1},L_{2}$ are parastrophic [^13] This terminology is borrowed from the quasigroup literature. In the Latin square literature this is sometimes referred to as “conjugate”, but we find the argument in Keedwell and Denes [35, pp. 15–16] to use the quasigroup nomenclature even in the setting of Latin squares compelling. if there is a permutation $\pi\in S_{3}$ (where $S_{3}$ is the symmetric group of degree $3$) such that, when viewed as sets of triples, we have $L_{2}=\{(x_{1^{\pi}},x_{2^{\pi}},x_{3^{\pi}}):(x_{1},x_{2},x_{3})\in L_{1}\}$; the induced map $L_{1}\to L_{2}$ is called a parastrophy. A main class isomorphism of Latin squares is the composition of a parastrophy and an isotopy; if there exists a main class isomorphism $L_{1}\to L_{2}$, we say they are main class isomorphic. [^14] Miller [43], and then Levet [37] following Miller, referred to this as “main class isotopy”; we have since found several modern textbooks on quasigroups and Latin squares that refer to this notion as one of “main class isomorphism”, “paratopy”, or “isostrophy” (sic). Keedwell and Denes [35, pp. 15–16] have a nice discussion of the terminology, as well as the history of its usage.
@@
For a given Latin square $L$ of order $n$, we associate a Latin square graph $G(L)$ that has $n^{2}$ vertices; one for each triple $(a,b,c)$ that satisfies $ab=c$. Two vertices $(a,b,c)$ and $(x,y,z)$ are adjacent in $G(L)$ precisely if $a=x$ or $b=y$ or $c=z$. Miller showed that two Latin square graphs $G_{1}$ and $G_{2}$ are isomorphic if and only if the corresponding Latin squares, $L_{1}$ and $L_{2}$, are main class isomorphic [43].
@@
A Latin square graph on $n^{2}$ vertices is a strongly regular graph with parameters $(n^{2},3(n-1),n,6)$. Conversely, a strongly regular graph with these same parameters $(n^{2},3(n-1),n,6)$ is called a pseudo-Latin square graph. Bruck showed that for $n>23$, a pseudo-Latin square graph is a Latin square graph [16].
Group Theory.
@@
For a standard reference, see [48]. All groups are assumed to be finite. For a group $G$, $d(G)$ denotes the minimum size of a generating set for $G$. The Frattini subgroup $\Phi(G)$ is the set of non-generators of $G$. If $G$ is a $p$-group, the Burnside Basis Theorem (see [48, Theorem 5.3.2]) provides that (i) $G=G^{p}[G,G]$, (ii) $G/\Phi(G)\cong(\mathbb{Z}/p\mathbb{Z})^{d(G)}$, and (iii) if $S$ generates $(\mathbb{Z}/p\mathbb{Z})^{d(G)}$, then any lift of $S$ generates $G$. A chief series of $G$ is an ascending chain $(N_{i})_{i=0}^{k}$ of normal subgroups of $G$, where $N_{0}=1$, $N_{k}=G$, and each $N_{i+1}/N_{i}$ ($i=0,\ldots,k-1$) is minimal normal in $G/N_{i}$. For $g,h\in G$, the commutator $g,h]:=ghg^{-1}h^{-1}$. The commutator subgroup $G,G]=\langle\{[g,h]:g,h\in G\}\rangle$.
Designs.
@@
Let $t\leq k\leq v$ and $\lambda$ be positive integers. A $(t,k,\lambda,v)$ design is an incidence structure $\mathcal{D}=(X,\mathcal{B},I)$, where $X$ is a set of $v$ points, $\mathcal{B}$ is a subset of $\binom{X}{k}$—whose elements are referred to as blocks—and such that each $t$-element subset of $X$ belongs to exactly $\lambda$ blocks. Now $I$ is the point-block incidence matrix, where $I_{x,B}=1$ precisely if the point $x$ belongs to the block $B$.
@@
For a design $\mathcal{D}=(X,\mathcal{B},I)$, we may define a block intersection graph (also known as a line graph) $G(V,E)$, where $V(G)=\mathcal{B}$ and two blocks $B_{1},B_{2}$ are adjacent in $G$ if and only if $B_{1}\cap B_{2}\neq\emptyset$. In the case of a Steiner $2$-design, the block-intersection graph is strongly regular. For Steiner triple systems, the block-intersection graphs are strongly regular with parameters (n(n1)/6,3(n3)/2,(n+3)/2,9)(n(n-1)/6,3(n-3)/2,(n+3)/2,9). Conversely, strongly regular graphs with the parameters (n(n1)/6,3(n3)/2,(n+3)/2,9)(n(n-1)/6,3(n-3)/2,(n+3)/2,9) are referred to as pseudo-STS graphs. Bose showed that pseudo-STS graphs with strictly more than $67$ vertices are in fact STS graphs [15].
Algorithmic Problems.
A multiplication table of a magma $G=\{g_{1},\dots,g_{n}\}$ of order $n$ is an array $M$ of length $n^{2}$ where each entry M[i+(j1)n]M[i+(j-1)n] for $i,j\in\{1,\dots,n\}$ contains the binary representation of $k$ such that $g_{i}g_{j}=g_{k}$. In the following all magmas are given as their multiplication tables.
@@
We will be primarily interested in Minimum Generating Set and Membership in the setting of (quasi)groups, with a few excursions to semigroups and magmas. Note that Membership for groups is known to be in $\mathsf{L}\cap\mathsf{quasiAC^{0}}$ [11, 47, 25]. Here the containment in $\mathsf{L}$ follows from the deep result by Reingold [47] that symmetric logspace (nondeterministic logspace where each transition is also allowed to be applied backward) coincides with $\mathsf{L}$.